OpenStudy (please.help.me):

√2/2 csc x − 1 = 0 Solve the equation. (Enter your answers as a comma-separated list. Use n as an integer constant. Enter your response in radians.)

OpenStudy (danjs):

This correct? \[\large \frac{ \sqrt{2} }{ 2 }*\csc(x)-1=0\]

OpenStudy (please.help.me):

Yes!

OpenStudy (mhchen):

\[\csc(x)=\frac{2}{\sqrt{2}}\] Right? And then \[sin(x)=\frac{\sqrt{2}}{{2}}\] Right?? Follow?

OpenStudy (mhchen):

And I think you should've memorized the unit circle for this problem. Right?

OpenStudy (please.help.me):

Yes!!!

OpenStudy (please.help.me):

What is the next step though?

OpenStudy (danjs):

you can take the inverse sine of both sides.. \[x=\sin^{-1}( \frac{ \sqrt{2} }{ 2 })=\] Or recall the unit circle. Are you familiar with the properties on the circle?

OpenStudy (please.help.me):

Are you asking what sine goes with one the unit circle?

OpenStudy (please.help.me):

\[\frac{ \pi }{ 4 }\]

OpenStudy (please.help.me):

7pi/4

OpenStudy (danjs):

remember, each point on the unit circle is (x,y)=(cos x, sin x)

OpenStudy (danjs):

The sine value is the y value of any point on the circle..

OpenStudy (please.help.me):

Oh! then it will go with 3pi/4. Right?

OpenStudy (danjs):

yes, this prob says sin(x)=root(2)/2, that occurs at 2 angles where the y coordinate value is root(2)/2 pi/4 and 3pi/4

OpenStudy (please.help.me):

Yes! I follow

OpenStudy (danjs):

That is the solutions if you limit the angle value to be 0 to 2pi. You can still keep going around the circle more times and get different angle values.. like pi/4 + 2pi

OpenStudy (danjs):

if you add a full revolution to both the angles, you are back at the same point on the circle

OpenStudy (please.help.me):

Yes, I understand!

OpenStudy (danjs):

So really the answer is those two angles, plus any multiple of 2pi pi/4 + 2pi*n 3pi/4 + 2pi*n for n=0,1,2,3...

OpenStudy (please.help.me):

So those are the final answers?

OpenStudy (please.help.me):

Also how do I know when to write + 2pi n or some problems the answer is just + 2 pi?

OpenStudy (danjs):

yes, and it depends on the solution. For this one, you got pi/4 and 3pi/4. a full revolution around the circle is 2pi. So if you are at pi/4 and add 2pi, you are back to the same spot with angle pi/4 + 2pi = 9pi/4

OpenStudy (danjs):

You can add as many full revolutions to both the angles and still be in the same point.. pi/4 + 2pi*n 3pi/4 + 2pi*n n=0,1,2...

OpenStudy (mhchen):